OpenStudy (anonymous):

A point M on a segment with endpoints X (1, −2) and Y (10, 3) partitions the segment in a 4:1 ratio. Find M. You must show all work to receive credit.

OpenStudy (anonymous):

So far I have Since I am given 4:1 I know to put it into a fraction which is 4/1. To find M I have to find the distance between the X coordinates and the y coordinates. From 1 to 10 you go 9 over to the right which makes it a positive 9. So now I will multiply 4/1 by 9. 4/1 X 9/1 = 36/1 = Now I will look at the Y coordinates which are -2 and 3. There is 5 spaces going up this is a positive 5. Now I will once again take 4/1 and multiply it by 5. 4/1 5/1 = 20/1 = 20 Now I have to find the exact coordinates of point M. but I know its incorrect iv graphed it and it makes no sense. Iv been trying to figure out this problem for an hour, anything helps.